OpenStudy (anonymous):

What is the value of the y variable in the solution to the following system of equations? 5x + 3y = 7 3x - 5y = -23

OpenStudy (anonymous):

\[5x + 3y = 7\] \[5x = 7 - 3y\] \[x= \frac{ 7 - 3y }{ 5 }\] Now sub in x into the second equation.

OpenStudy (anonymous):

Thanks!!!

OpenStudy (anonymous):

But I still don't get it.

OpenStudy (anonymous):

Remember to put the new found y value back into the first one to get the numerical value for x

OpenStudy (anonymous):

Alright so basically what you've done so far is isolate x in the first equation. Now you use that as your "x" value in the second equation. When you do that you'll end up with a y value. When you get this y value plug it back into the first equation and you'll get the x value. Now you have both the x and y. Lemme know if you need help

OpenStudy (anonymous):

Ohh ok now I get it. x=-19 and y=24 is that correct?? @SA5UK3

OpenStudy (anonymous):

@SA5UK3

OpenStudy (anonymous):

Umm I don't think so. Let me show you what I did

OpenStudy (anonymous):

\[3(\frac{ 7-3y }{ 5 }) - 5y = -23\] \[\frac{ 21-9y }{ 5 } -5y = -23\] \[\frac{ 21-9y-25y }{ 5 } = -23\] \[21-9y-25=-113\] \[21-34y=-115\] \[-34y = -136\] \[y = \frac{ -136 }{ -34 }\] \[y = 4\]

OpenStudy (anonymous):

woah that's a lot of math. THANK YOU SO MUCH!!!!!!!!!!!!!!!!!!!! @SA5UK3 Your medal will be with you shortly ^-^

OpenStudy (anonymous):

Haha, its not really a lot. I just tried to show every step. So now that you have your y value put it back into the first equation and you'll get your x.